# Integrated Optimal Control for Fast Charging and Active Thermal Management of Lithium-Ion Batteries in Extreme Ambient Temperatures

### v1

@@ V EKF-Based Output-Feedback Model Predictive Control
The previous section studies the MPC of battery charging based on the assumption that the states are fully measurable. This assumption might not hold in real-world applications. Hence, this section proposes an output-feedback MPC based on Extended Kalman Filter (EKF) [35, Chapter 8.2]. The EKF takes in system output measurements and generates state estimates, which are given to a revised MPC controller to design control. This section presents a way to rewrite the proposed thermal-NDC model for the ease of EKF, proposes an EKF-based output-feedback MPC, and demonstrates the effectiveness of this algorithm under extreme ambient temperatures by numerical simulations.
@@ V-A System Reformulation & Proposed Algorithm
The output $V$ is a function of both states and inputs, as shown in (1f), thus EKF is not applicable for this output form. This subsection first rewrites the original system dynamics (10) similarly as [28] such that its output is a function of states. To reformulate the system dynamics, new state, control, and output for the continuous-time system dynamics are defined as follows,
@@
where $u_{1}\in\mathbb{R}$ is a new control input such that the charging current has the following dynamics:
@@
Combining the previous dynamics (II) - (8) with the current dynamics (24), one obtains the following dynamics in continuous-time:
@@
The EKF for the battery system (26) is summarized in Algorithm 2. Denote $\boldsymbol{F}_{i}\triangleq\frac{\partial\boldsymbol{f}_{\mathrm{d}}(\boldsymbol{x},\boldsymbol{u})}{\partial\boldsymbol{x}}|_{\hat{\boldsymbol{x}}_{i-1|i-1},\boldsymbol{u}_{i-1}}$ and $\boldsymbol{H}_{i}\triangleq\frac{\partial\boldsymbol{h}(\boldsymbol{x})}{\partial\boldsymbol{x}}|_{\hat{\boldsymbol{x}}_{i-1|i-1}}.$ At each time instance $t_{i}$, the system output $\boldsymbol{y}_{i}$ is observed. Then with the previous state estimate $\hat{\boldsymbol{x}}_{i-1|i-1}$, covariance estimate $\boldsymbol{P}_{i-1|i-1}$, and control input $\boldsymbol{u}_{i-1}$, the EKF updates the present estimates $\hat{\boldsymbol{x}}_{i|i}$ and $\boldsymbol{P}_{i|i}$. Note that the estimated SOC at time $t_{i}$ is given by its definition (3), i.e. $\hat{\mathrm{SOC}}(t_{i})=(C_{\mathrm{b}}\hat{V}_{\mathrm{b}}(t_{i})+C_{\mathrm{s}}\hat{V}_{\mathrm{s}}(t_{i}))/(C_{\mathrm{b}}+C_{\mathrm{s}})$.
@@
2       $\hat{\boldsymbol{x}}_{i|i-1}\leftarrow\boldsymbol{f}_{\mathrm{d}}(\hat{\boldsymbol{x}}_{i-1|i-1},\boldsymbol{u}_{i-1})$ $\boldsymbol{P}_{i|i-1}=\boldsymbol{F}_{i}\boldsymbol{P}_{i-1|i-1}\boldsymbol{F}_{i}^{\top}+\boldsymbol{Q}_{i}$ $\tilde{\boldsymbol{y}}_{i}\leftarrow\boldsymbol{y}_{i}-\boldsymbol{h}(\hat{\boldsymbol{x}}_{i|i-1})$ $\boldsymbol{S}_{i}\leftarrow\boldsymbol{H}_{i}\boldsymbol{P}_{i|i-1}\boldsymbol{H}_{k}^{\top}+\boldsymbol{R}_{i}$ $\boldsymbol{K}_{i}\leftarrow\boldsymbol{P}_{i|i-1}\boldsymbol{H}_{i}^{\top}\boldsymbol{S}_{i}^{-1}$ $\hat{\boldsymbol{x}}_{i|i}=\hat{\boldsymbol{x}}_{i|i-1}+\boldsymbol{K}_{i}\tilde{\boldsymbol{y}}_{i}$ $\boldsymbol{P}_{i|i}=(\boldsymbol{I}_{5}-\boldsymbol{K}_{i}\boldsymbol{H}_{i})\boldsymbol{P}_{i|i-1}$ return $\hat{\boldsymbol{x}}_{i|i},\boldsymbol{P}_{i|i}$
@@
1 initialize $\hat{\boldsymbol{x}}_{i-1|i-1}$ and $\boldsymbol{P}_{i-1|i-1}$ while $\hat{\mathrm{SoC}}_{i}<\mathrm{SoC}_{\mathrm{r}}$ do
2       observe $\boldsymbol{y}_{i}$ and $T_{\mathrm{amb},i}$ at time $t_{i}$ $\hat{\boldsymbol{x}}_{i|i},\boldsymbol{P}_{i|i}\leftarrow$ estimate($\hat{\boldsymbol{x}}_{i-1|i-1},\boldsymbol{P}_{i-1|i-1},\boldsymbol{u}_{i-1},\boldsymbol{y}_{i}$) obtain $\hat{\mathrm{SoC}}_{i}$ given $\hat{\boldsymbol{x}}_{i|i}$ if $t_{i}\ \%\ \Delta_{\mathrm{p}}==0$ then
@@ V-B Numerical Simulations
The covariance matrix $\boldsymbol{Q}_{i}$ for the process noise is set based on the relative modeling error or process uncertainty. For $V_{\mathrm{b}}$ and $V_{\mathrm{s}}$, the dominant source of error is the capacitor $C_{\mathrm{b}}$, as described in (1f). To account for a relative error of $\pm$ 7.5% with a probability of 99.7%, the variance for $V_{\mathrm{b}}$ in discrete-time should be $(2.5\%\Delta_{\mathrm{s}}\frac{-1}{C_{\mathrm{b}}R_{\mathrm{b}}})^{2}=1.73\times 10^{-8}$. The variance for $C_{\mathrm{s}}$ is also $1.73\times 10^{-8}$ for conservatism. For $T_{\mathrm{core}}$ and $T_{\mathrm{surf}}$, according to (6) and the measurability of $T_{\mathrm{surf}}$, the dominant source of error is $T_{\mathrm{core}}$. To cover a relative error of $\pm$ 7.5% with a probability of 99.7%, the variance for $T_{\mathrm{core}}$ should be $(2.5\%\Delta_{\mathrm{s}}\frac{-1}{R_{\mathrm{core}}C_{\mathrm{core}}})^{2}=2.44\times 10^{-8}$. For $T_{\mathrm{surf}}$, to cover a relative error of $\pm$ 0.3%, the variance should be $(0.1\%\Delta_{\mathrm{s}}(\frac{-1}{R_{\mathrm{surf}}C_{\mathrm{surf}}}+\frac{-1}{R_{\mathrm{core}}C_{\mathrm{surf}}}))^{2}=1.54\times 10^{-9}$. Since the current dynamics (24) is manually constructed to enable EKF, there is no modeling error or uncertainty on $I$, thus the variance for $I$ is 0. Therefore, the process covariance matrix is $\boldsymbol{Q}_{i}=\text{diag}(1.73\times 10^{-8},\ 1.73\times 10^{-8},\ 2.44\times 10^{-8},\ 1.54\times 10^{-9},\ 0)$.
@@
    High Ambient Temperature: $T_{\mathrm{amb}}=70$ ${}^{\circ}\text{C}$; $\boldsymbol{x}_{0}=\begin{bmatrix}0.1\text{ V}&0.1\text{ V}&50\,^{\circ}\text{C}&70\,^{\circ}\text{C}&0\text{ A}\end{bmatrix}^{\top}$;
@@
The initial estimates on the unmeasurable states $V_{\mathrm{b}}$ and $T_{\mathrm{core}}$ are randomly initialized within $[V_{\mathrm{b}}(0)-0.1\mathrm{V},V_{\mathrm{b}}(0)+0.1\mathrm{V}]$ and $[T_{\mathrm{core}}(0)-5\ ^{\circ}\text{C},T_{\mathrm{core}}(0)+5\ ^{\circ}\text{C}]$, respectively. Note that the initial estimate error on $V_{\mathrm{b}}$ is relatively large since $V_{\mathrm{b}}$ varies between $[0\mathrm{V},1\mathrm{V}]$. The initial estimates of the measurable states are given by the initial measurements. For each $T_{\mathrm{amb}}$, the proposed EKF-based output-feedback MPC runs 20 trials with random initial estimates to verify its effectiveness.
@@
${}^{\dagger}$ $T_{\mathrm{amb}}$ mean $\pm$ std ${}^{\ddagger}$
$V_{\mathrm{b}}$ 25 ${}^{\circ}$ $0.0012\pm 0.0064$ 0.0003, 0.0007, 0.0011
$V_{\mathrm{s}}$ 25 ${}^{\circ}$ $0.0008\pm 0.0018$ 0.0003, 0.0006, 0.0009
$T_{\mathrm{core}}$ 25 ${}^{\circ}$ $0.0172\pm 0.1051$ 0.0013, 0.0030, 0.0092
$\mathrm{SoC}$ 25 ${}^{\circ}$ $0.11\%\pm 0.59\%$ 0.03%, 0.06%, 0.11%
$V_{\mathrm{b}}$ 70 ${}^{\circ}$ $0.0012\pm 0.0065$ 0.0003, 0.0007, 0.0011
$V_{\mathrm{s}}$ 70 ${}^{\circ}$ $0.0008\pm 0.0018$ 0.0003, 0.0006, 0.0009
$T_{\mathrm{core}}$ 70 ${}^{\circ}$ $0.0092\pm 0.1070$ 0.0011, 0.0023, 0.0047
$\mathrm{SoC}$ 70 ${}^{\circ}$ $0.11\%\pm 0.60\%$ 0.03%, 0.06%, 0.11%
$V_{\mathrm{b}}$25 ${}^{\circ}$ $0.0012\pm 0.0067$ 0.0003, 0.0006, 0.0011
$V_{\mathrm{s}}$25 ${}^{\circ}$ $0.0008\pm 0.0018$ 0.0003, 0.0005, 0.0009
$T_{\mathrm{core}}$25 ${}^{\circ}$ $0.0174\pm 0.1087$ 0.0018, 0.0056, 0.0186
$\mathrm{SoC}$25 ${}^{\circ}$ $0.11\%\pm 0.62\%$ 0.03%, 0.06%, 0.11%
@@
    error unit for $V_{\mathrm{b}}$, $V_{\mathrm{s}}$, $T_{\mathrm{core}}$, and $\mathrm{SoC}$ are V, V, ${}^{\circ}$, and %, respectively
@@
The charging performance for all cases is summarized in Table IX. Comparing with the performance of P1 in Table III - Table V, both the computational time and the charging time from Table IX are longer than those of the state-feedback MPC. One contributing factor is the conservative revision of the Li-ion concentration gradient constraint (28), which includes a safety margin to account for estimation error and measurement noise. To illustrate this, consider the trajectory of $V_{\mathrm{s}}(t)-V_{\mathrm{b}}(t)$ vs $\beta_{1}\mathrm{SoC}(t)+\beta_{2}$ with $T_{\mathrm{amb}}=70$$^{\circ}$, as shown in Fig. 9. In Fig. 3, the same trajectory of $V_{\mathrm{s}}(t)-V_{\mathrm{b}}(t)$ without any noise or uncertainty approaches the boundary of $\beta_{1}\mathrm{SoC}(t)+\beta_{2}$ as SoC exceeds about 70%. However, when using output-feedback MPC with noise and uncertainty, the actual $V_{\mathrm{s}}(t)-V_{\mathrm{b}}(t)$ could exceed this boundary. Therefore, the 5% safety margin in (28) prevents constraint violation, but it also slows down the charging process due to the reduced concentration gradient.
@@
VI Conclusion and Future Work
This paper explores integrated control strategies for fast charging and active thermal management of LiBs under extreme ambient temperatures. Firstly, a control-oriented thermal-NDC model is introduced to describe the electrical and thermal dynamics of LiBs, including the impact from both an active thermal source and ambient temperature. Secondly, a state-feedback MPC algorithm is developed to concurrently perform fast charging and active thermal management.
@@ VI Conclusion and Future Work
Thirdly, numerical experiments are conducted to validate the proposed state-feedback MPC algorithm under extreme ambient temperatures. These experiments demonstrate that heating the battery can enhance fast charging, which is consistent with the literature. By explicitly incorporating a higher desired battery core temperature into the MPC’s objective function, the algorithm achieves the fastest charging speed. However, with a suitable prediction horizon, an objective function without explicit heating can achieve a balanced performance across various indices. Finally, the proposed thermal-NDC model is rewritten properly to enable an EKF design for battery state estimation. An EKF-based output-feedback MPC algorithm is then proposed and validated through numerical experiments under extreme ambient temperatures.
The proposed thermal-NDC model is built upon the battery charging and discharging data, utilizing a maximum current of 1.5 C. As noted by [8, 19], applications such as eVTOL aircraft and EVs require high C-rate charging, which leads to significant heat generation and poses potentially a bigger concern on system safety. One avenue for future work is to develop battery model suitable for higher charging currents. Another future direction involves developing integrated charging and thermal management strategies for high C-rate scenarios. This may also require detailed modeling of actuators in the heating and cooling system.

### v3

@@ V EKF-Based Output-Feedback Model Predictive Control
The previous section studies the MPC of battery charging based on the assumption that the states are fully measurable. This assumption might not hold in real-world applications. Hence, this section proposes an output-feedback MPC based on Extended Kalman Filter (EKF) [49, Chapter 8.2]. The EKF takes in system output measurements and generates state estimates, which are given to a revised MPC controller to design control. This section presents a way to rewrite the proposed thermal-NDC model for the ease of EKF, proposes an EKF-based output-feedback MPC, and demonstrates the effectiveness of this algorithm under extreme ambient temperatures by numerical simulations.
@@ V-A System Reformulation & Proposed Algorithm
The output $V$ is a function of both states and inputs, as shown in (1b), thus EKF is not applicable for this output form. This subsection first rewrites the original system dynamics (10) similarly as [28] such that its output is a function of states. To reformulate the system dynamics, new state, control, and output for the continuous-time system dynamics are defined as follows,
@@
where $u_{1}\in\mathbb{R}$ is a new control input such that the charging current has the following fictitious dynamics:
@@
Combining the previous dynamics (1) - (8) with the current dynamics (24), one obtains the following dynamics in continuous-time:
@@
The EKF for the battery system (26) is summarized in Algorithm 2. Denote $\boldsymbol{F}_{i}\triangleq\frac{\partial\boldsymbol{f}_{\mathrm{d}}(\boldsymbol{x},\boldsymbol{u})}{\partial\boldsymbol{x}}|_{\hat{\boldsymbol{x}}_{i-1|i-1},\boldsymbol{u}_{i-1}}$ and $\boldsymbol{H}_{i}\triangleq\frac{\partial\boldsymbol{h}(\boldsymbol{x})}{\partial\boldsymbol{x}}|_{\hat{\boldsymbol{x}}_{i-1|i-1}}.$ At each time instance $t_{i}$, the system output $\boldsymbol{y}_{i}$ is observed. Then with the previous state estimate $\hat{\boldsymbol{x}}_{i-1|i-1}$, covariance estimate $\boldsymbol{P}_{i-1|i-1}$, and control input $\boldsymbol{u}_{i-1}$, the EKF updates the present estimates $\hat{\boldsymbol{x}}_{i|i}$ and $\boldsymbol{P}_{i|i}$. Note that the estimated SoC at time $t_{i}$ is given by its definition (3), i.e., $\hat{\mathrm{SoC}}(t_{i})=(C_{\mathrm{b}}\hat{V}_{\mathrm{b}}(t_{i})+C_{\mathrm{s}}\hat{V}_{\mathrm{s}}(t_{i}))/(C_{\mathrm{b}}\overline{V}_{\mathrm{b}}+C_{\mathrm{s}}\overline{V}_{\mathrm{s}})$.
@@
2       $\hat{\boldsymbol{x}}_{i|i-1}\leftarrow\boldsymbol{f}_{\mathrm{d}}(\hat{\boldsymbol{x}}_{i-1|i-1},\boldsymbol{u}_{i-1})$
3       $\boldsymbol{P}_{i|i-1}=\boldsymbol{F}_{i}\boldsymbol{P}_{i-1|i-1}\boldsymbol{F}_{i}^{\top}+\boldsymbol{Q}_{i}$
4       $\tilde{\boldsymbol{y}}_{i}\leftarrow\boldsymbol{y}_{i}-\boldsymbol{h}(\hat{\boldsymbol{x}}_{i|i-1})$
5       $\boldsymbol{S}_{i}\leftarrow\boldsymbol{H}_{i}\boldsymbol{P}_{i|i-1}\boldsymbol{H}_{k}^{\top}+\boldsymbol{R}_{i}$
6       $\boldsymbol{K}_{i}\leftarrow\boldsymbol{P}_{i|i-1}\boldsymbol{H}_{i}^{\top}\boldsymbol{S}_{i}^{-1}$
7       $\hat{\boldsymbol{x}}_{i|i}=\hat{\boldsymbol{x}}_{i|i-1}+\boldsymbol{K}_{i}\tilde{\boldsymbol{y}}_{i}$
8       $\boldsymbol{P}_{i|i}=(\boldsymbol{I}_{5}-\boldsymbol{K}_{i}\boldsymbol{H}_{i})\boldsymbol{P}_{i|i-1}$
9       return $\hat{\boldsymbol{x}}_{i|i},\boldsymbol{P}_{i|i}$
10
@@
1 initialize $\hat{\boldsymbol{x}}_{i-1|i-1}$ and $\boldsymbol{P}_{i-1|i-1}$
2 while $\hat{\mathrm{SoC}}_{i}<\mathrm{SoC}_{\mathrm{r}}$ do
3       observe $\boldsymbol{y}_{i}$ and $T_{\mathrm{amb},i}$ at time $t_{i}$
4       $\hat{\boldsymbol{x}}_{i|i},\boldsymbol{P}_{i|i}\leftarrow$ estimate($\hat{\boldsymbol{x}}_{i-1|i-1},\boldsymbol{P}_{i-1|i-1},\boldsymbol{u}_{i-1},\boldsymbol{y}_{i}$)
5       obtain $\hat{\mathrm{SoC}}_{i}$ given $\hat{\boldsymbol{x}}_{i|i}$
6
7      if $t_{i}\ \%\ \Delta_{\mathrm{p}}==0$ then
8             $k\leftarrow k+1$
@@ V-B Numerical Simulations
The covariance matrix $\boldsymbol{Q}_{i}$ for the process noise is set based on the relative modeling error or process uncertainty. For $V_{\mathrm{b}}$ and $V_{\mathrm{s}}$, the dominant source of error is the capacitor $C_{\mathrm{b}}$, as described in (1b). To account for a relative error of $\pm$ 7.5% with a probability of 99.7%, the variance for $V_{\mathrm{b}}$ in discrete-time should be $(2.5\%\Delta_{\mathrm{s}}\frac{-1}{C_{\mathrm{b}}R_{\mathrm{b}}})^{2}=1.73\times 10^{-8}$. The variance for $C_{\mathrm{s}}$ is also $1.73\times 10^{-8}$ for conservatism. For $T_{\mathrm{core}}$ and $T_{\mathrm{surf}}$, according to (6) and the measurability of $T_{\mathrm{surf}}$, the dominant source of error is $T_{\mathrm{core}}$. To cover a relative error of $\pm$ 7.5% with a probability of 99.7%, the variance for $T_{\mathrm{core}}$ should be $(2.5\%\Delta_{\mathrm{s}}\frac{-1}{R_{\mathrm{core}}C_{\mathrm{core}}})^{2}=2.44\times 10^{-8}$. For $T_{\mathrm{surf}}$, to cover a relative error of $\pm$ 0.3%, the variance should be $(0.1\%\Delta_{\mathrm{s}}(\frac{-1}{R_{\mathrm{surf}}C_{\mathrm{surf}}}+\frac{-1}{R_{\mathrm{core}}C_{\mathrm{surf}}}))^{2}=1.54\times 10^{-9}$. Since the current dynamics (24) is manually constructed to enable EKF, there is no modeling error or uncertainty on $I$, thus the variance for $I$ is 0. Therefore, the process covariance matrix is $\boldsymbol{Q}_{i}=\text{diag}(1.73\times 10^{-8},\ 1.73\times 10^{-8},\ 2.44\times 10^{-8},\ 1.54\times 10^{-9},\ 0)$.
@@
    High Ambient Temperature: $T_{\mathrm{amb}}=70$ ${}^{\circ}\text{C}$, $\boldsymbol{x}_{0}=\begin{bmatrix}0.1\text{ V}&0.1\text{ V}&50\,^{\circ}\text{C}&70\,^{\circ}\text{C}&0\text{ A}\end{bmatrix}^{\top}$;
@@
[…]
@@
State $T_{\mathrm{amb}}$ mean $\pm$ std quartiles
$V_{\mathrm{b}}$ 25 C $0.0012\pm 0.0064$ 0.0003, 0.0007, 0.0011
$V_{\mathrm{s}}$ 25 C $0.0008\pm 0.0018$ 0.0003, 0.0006, 0.0009
$T_{\mathrm{core}}$ 25 C $0.0172\pm 0.1051$ 0.0013, 0.0030, 0.0092
$\mathrm{SoC}$ 25 C $0.11\%\pm 0.59\%$ 0.03%, 0.06%, 0.11%
$V_{\mathrm{b}}$ 70 C $0.0012\pm 0.0065$ 0.0003, 0.0007, 0.0011
$V_{\mathrm{s}}$ 70 C $0.0008\pm 0.0018$ 0.0003, 0.0006, 0.0009
$T_{\mathrm{core}}$ 70 C $0.0092\pm 0.1070$ 0.0011, 0.0023, 0.0047
$\mathrm{SoC}$ 70 C $0.11\%\pm 0.60\%$ 0.03%, 0.06%, 0.11%
$V_{\mathrm{b}}$25 C $0.0012\pm 0.0067$ 0.0003, 0.0006, 0.0011
$V_{\mathrm{s}}$25 C $0.0008\pm 0.0018$ 0.0003, 0.0005, 0.0009
$T_{\mathrm{core}}$25 C $0.0174\pm 0.1087$ 0.0018, 0.0056, 0.0186
$\mathrm{SoC}$25 C $0.11\%\pm 0.62\%$ 0.03%, 0.06%, 0.11%
@@
    error unit for $V_{\mathrm{b}}$, $V_{\mathrm{s}}$, $T_{\mathrm{core}}$, and $\mathrm{SoC}$ are V, V, K, and %, respectively
@@
The charging performance for all cases is summarized in Table IX. Comparing with the performance of P1 in Table III - Table V, both the computational time and the charging time from Table IX are longer than those of the state-feedback MPC. One contributing factor is the conservative revision of the Li-ion concentration gradient constraint (28), which includes a safety margin to account for estimation error and measurement noise. To illustrate this, consider the trajectory of $V_{\mathrm{s}}(t)-V_{\mathrm{b}}(t)$ vs $\beta_{1}\mathrm{SoC}(t)+\beta_{2}$ with $T_{\mathrm{amb}}=70$ C, as shown in Fig. 9. In Fig. 3, the same trajectory of $V_{\mathrm{s}}(t)-V_{\mathrm{b}}(t)$ without any noise or uncertainty approaches the boundary of $\beta_{1}\mathrm{SoC}(t)+\beta_{2}$ as SoC exceeds about 70%. However, when using output-feedback MPC with noise and uncertainty, the actual $V_{\mathrm{s}}(t)-V_{\mathrm{b}}(t)$ could exceed this boundary. Therefore, the 5% safety margin in (28) prevents constraint violation, but it also slows down the charging process due to the reduced concentration gradient.
@@
VI Conclusion and Discussion
@@
This paper explores integrated control strategies for optimal fast charging and active thermal management of LiBs under extreme ambient temperatures, aiming to balance charging speed and battery health. First, a control-oriented thermal-NDC model is introduced to describe the electrical and thermal dynamics of LiBs, including the impact from both an active thermal source and ambient temperature. Second, a state-feedback MPC algorithm is developed to perform optimal fast charging and active thermal management concurrently. Third, numerical experiments are conducted to validate the proposed state-feedback MPC algorithm under extreme ambient temperatures. These experiments demonstrate that heating the battery can enhance fast charging, which is consistent with the literature. By explicitly incorporating a higher desired battery core temperature into the MPC’s objective function, the algorithm achieves faster charging speed. However, with a suitable prediction horizon, an objective function without explicit heating can achieve a balanced performance across various indices. Finally, the proposed thermal-NDC model is rewritten properly to enable an EKF design for battery state estimation. An EKF-based output-feedback MPC algorithm is then proposed and validated through numerical experiments under extreme ambient temperatures.
The limitations of this paper can be summarized from three key perspectives. First, the proposed thermal-NDC model is validated for charging and discharging currents up to 1.5C. This limitation arises because the model is based on experimental data collected at a maximum current of 1.5C, limiting its applicability. As highlighted by [8, 19], applications such as eVTOL aircraft and EVs often require high C-rate charging, which generates significant heat and raises concerns about system safety. However, the methodology presented in this paper could be readily adapted to higher C-rate charging if a more sophisticated model, suitable for high C-rates, is used. Therefore, a future work direction is to develop a battery model and corresponding control strategy capable of handling higher currents, such as up to 4-6C.
Second, the lumped thermal model used in this paper may lose some fidelity, as higher-fidelity thermal dynamics are described by PDEs [38]. Nevertheless, lumped thermal models, governed by ODEs, are frequently used in the literature [17, 18, 24, 36, 42, 44, 45] due to the computational feasibility for real-time model-based control. Some studies utilize two-point or multiple-point lumped models to capture the thermal dynamics of battery packs [42, 44]. Although the literature [38] and experimental data [17] suggest that lumped thermal models remain valid for control purposes, this paper acknowledges the modeling mismatch between lumped thermal models and actual thermal dynamics. This discrepancy highlights the need to design robust control strategies to address modeling errors and process uncertainties in future work.
Third, the dynamics modeling of the heating/cooling system is omitted in this paper, simplifying the active thermal power as (8). Although modeling the heating/cooling system with actuator dynamics (e.g., heat pump, condenser, evaporator, fan, etc.) provides valuable insights into system behavior and benefits model-based thermal management [53, 54, 55], the additional complexity would detract from the primary focus of this paper. The advantage of using (8) is that it simplifies the modeling and control of the battery cell and thermal system by hierarchically separating them. This approach allows the proposed algorithm to plan and determine the desired active thermal power for the battery cell. Subsequently, a separate control strategy can regulate the thermal system’s actuators to achieve this desired thermal power. Another benefit of using (8) is modularization. As noted in [42], the lumped model for a battery cell can be extended as a multi-point lumped model for a battery pack, where one node of the battery cells receives active thermal power through cold plates. The above justifications make the lumped model for thermal dynamics a reasonable choice for battery cell electrothermal dynamics and model-based control strategies. Regarding the typical range for the value of $\eta_{\mathrm{act}}$ in (8), it is difficult to specify one because this efficiency depends on various factors, such as the configuration of the heating/cooling system, the topology of the battery pack, actuator types, etc. Furthermore, as noted in [53, 54, 55], this efficiency is typically a state variable of the thermal system, influenced by fan speed, pump speed, coolant temperature, and other factors. Therefore, if a constant value for $\eta_{\mathrm{act}}$ cannot be predetermined, $\dot{Q}_{\mathrm{act}}(t)$ can be made as a decision variable. The proposed strategy can then determine a trajectory for the desired net active thermal power. A separate strategy from the heating/cooling system can then be employed to track this desired power trajectory as a reference trajectory. Nevertheless, this limitation underscores the need for future work to model the thermal dynamics and the heating/cooling dynamics (including efficiency) for a battery pack.